Circular No. 5879

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993ac IN ANONYMOUS GALAXY Jean Mueller reports her discovery of a supernova (mag ∼ 18) on a

IIIa-J plate taken by herself on Oct. 13 UT with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. The new object is located 4′′ west and 31′′.6 north of the nucleus of a galaxy at α = 5^{h}41^{m}.5, δ = +63^{o}20′ (equinox 1950.0). A spectrogram obtained by S. G. Djorgovski and D. Thompson on Oct. 14 with the 5-m Hale telescope (+ double spectrograph) shows broad Balmer emission lines with P-Cyg profiles and a very blue continuum.

SNR G10.0–0.3 S. Kulkarni, California Institute of Technology; D. Frail, National Ra-

dio Astronomy Observatory; N. Kassim, Naval Research Laboratory; T. Murakami, Institute for Space and Astronautical Science; and G. Vasisht, California Institute of Technology, write: “We have obtained multi-band images from data obtained at the Very Large Array on Sept. 13 of the supernova remnant (SNR) G10.0–0.3, which we had earlier claimed to be associated with the soft gamma repeater SGR 1806–20 (Kulkarni and Frail 1993, Nature 365, 33). The new images show a compact nebula superposed on an extended plateau emission. The morphology is suggestive of a plerion (i.e., a pulsar-powered SNR). The centroid of the compact nebula, the most likely position of the putative pulsar, is at α = 18^{h}05^{m}41^{s}.76, δ = −20^{o}25′13′′ (equinox 1950.0). Since plerions are quite rare, the association of SNR G10.0–0.3 with SGR 1806–20 should now be taken seriously. We speculate that the seat of the SGR activity lies in this compact nebula. Given that SGR 1806–20 is currently active (IAUC 5875), we urge multiwavelength — especially infrared (to minimize extinction) and x-ray — observations towards this position.”

GRO J0422+32 P. Zhao, P. Callanan, M. Garcia, and J. McClintock, Smithsonian

Astrophysical Observatory, communicate: “GRO J0422+32 has declined rapidly since Sept. 20. CCD photometry obtained at the F. L. Whipple Observatory 1.2-m telescope gives the following R magnitudes: Sept. 21.47 UT, 17.75 ± 0.02; 22.47, 17.84 ± 0.02; 23.49, 18.19 ± 0.02; Oct. 3.42, 18.75 ± 0.08; 8.451, 18.72 ± 0.02; 8.464, 18.71 ± 0.02.”

1993 October 14 Daniel W. E. Green